OpenStudy (anonymous):

Find the approximate volume of the sugar cone if the radius is 4 cm and the height is 12 cm. il solve like for the volume of a cylinder correct but divided by 2 ?

OpenStudy (anonymous):

volume= 1/3 (area of the base)(height)

OpenStudy (diyadiya):

\[Voluume \ of \ a \ cylider = \frac{1}{3} \times \pi \times (radius)^2 \times \ height\]Why do you have to divide by 2?

OpenStudy (anonymous):

1/3(4 cm)2(12 cm) = 1/3(16 cm2)(12 cm) 201.06 cm3 but i also found that the surface area is 402.12/2 gives you the same and thats what i did

OpenStudy (diyadiya):

You forgot π

OpenStudy (diyadiya):

Value of π is 3.14

OpenStudy (anonymous):

alright thank you

OpenStudy (diyadiya):

1/3 x (3.14) x (16) x (12) = 1/3 x 602.88 = 200.96 200.96 is approximately equal to 402.12/2 -You're Welcome :)